The time until recharge for a battery in a laptop computer under common conditions is normally distributed with mean of 200 minutes and a standard deviation of 60 minutes. What is the probability that a battery lasts more than 3 hours?
Answer #1

SOLUTION-

TIME FOR A BATTERY~NORMAL(200 MINUTES, 60 MINUTES)

SO,

3 HOURS= 180 MINUTES

P(X>180) = 1 - P( )

= 1 - P()

= 1 - P(Z -0.33)

= 1 - ; WHERE  Z IS A STANDARD NORMAL VARIABLE

= 1 - 1 + ; AS

=

= 0.6293

**REMARK**- IN CASE OF DOUBT, COMMENT BELOW AND DO LIKE THE SOLUTION IF YOU LIKED IT.